OpenStudy (anonymous):

Please how you got your answer. Edwin made two signboards in different shapes but with the same area. The first one is triangular with the dimensions shown below. The second signboard is rectangular and it is the same length as the base of the triangular signboard. Which is closest to the width in inches of the rectangular signboard? A. 10 B. 15 C. 20 D. 30

OpenStudy (anonymous):

Picture:

OpenStudy (anonymous):

the problem is to find the width of the rectangle?

OpenStudy (anonymous):

use this formula: area of the triangle = area of the rectangle 1/2(b)(h) = (b)(h) 1/2(42)(28) =(28)(h) find each which represent the width of the rectangle. . can you do it now?

OpenStudy (anonymous):

i mean find h. .

OpenStudy (anonymous):

I think so...

OpenStudy (anonymous):

Does h equal to 21?

OpenStudy (anonymous):

yeah h is 21. . but the question is "what is the nearest measure of the width," right? then, what is the number on the options closest to to 21.

OpenStudy (anonymous):

20? Cause 21 can be rounded down to 20..

OpenStudy (anonymous):

yeah.. correct

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

glad to help. .